# Functional differential equations driven by càdlàg rough paths

### v1

@@ 1. Introduction
Organization of the paper: In Section 2 we provide existence, uniqueness and continuity results for rough functional differential equations. In Section 3 we prove that various classes of rough differential equations are covered by the presented results on rough functional differential equations. In Section 4 we establish the existence of the Itô rough path lift of delayed martingales and discuss applications to stochastic differential equations with delay. Appendix A contains some auxiliary estimates for rough integrals.
Acknowledgments: A. P. Kwossek and D. J. Prömel gratefully acknowledge financial support by the Baden-Württemberg Stiftung.
@@ 2. Existence, uniqueness and continuity
Let us consider the rough functional differential equation (RFDE)
where $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ is a càdlàg $p$-rough path for $p\in(2,3)$, $(y,y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ is a given controlled path with respect to $X$ and $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ is a non-anticipative functional, i.e.
  • (i)
  • (ii)
for every $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$. The integral in (2.1) is defined as a (forward) rough integral, see (2.2) below for its definition. Note that the RFDE (1.1) can be re-written in the form of (2.1), using a standard time-extension of the driving rough path.
Before studying the RFDE (2.1), we will recall the necessary definitions and some essentials from the theory of càdlàg rough paths, as introduced by Friz and Shekhar [FS17] and Friz and Zhang [FZ18].
@@ 2.1. Essentials on rough path theory
Throughout, let $T>0$ be a fixed finite time horizon. Let $\Delta_{T}:=\{(s,t)\in[0,T]^{2}\,:\,s\leq t\}$ be the standard $2$-simplex. A function $w\colon\Delta_{T}\to[0,\infty)$ is called a control function if it is superadditive, in the sense that $w(s,u)+w(u,t)\leq w(s,t)$ for all $0\leq s\leq u\leq t\leq T$.
Whenever $(B,\|\cdot\|)$ is a normed space and $f,g\colon B\to\mathbb{R}$ are two functions on $B$, we shall write $f\lesssim g$ or $f\leq Cg$ to mean that there exists a constant $C>0$ such that $f(x)\leq Cg(x)$ for all $x\in B$. The constant $C$ may depend on the normed space, e.g. through its dimension or regularity parameters, and, if we want to emphasize the dependence of the constant $C$ on some particular variables, say $\alpha_{1},\ldots,\alpha_{n}$, then we will write $C=C_{\alpha_{1},\ldots,\alpha_{n}}$. Unless otherwise stated, the dependence of the implicit constants on the variables is locally bounded; that is, if $\alpha_{1}\in A_{1},\ldots,\alpha_{n}\in A_{n}$, where $A_{1},\ldots,A_{n}$ are compact subsets of the range of $\alpha_{1},\ldots,\alpha_{n}$ respectively, then we have that $\sup_{\alpha_{1}\in A_{1},\ldots,\alpha_{n}\in A_{n}}C_{\alpha_{1},\ldots,\alpha_{n}}<\infty$.
@@
for $(s,t)\in\Delta_{T}$. We point out that, by definition,
@@
exists (in the classical mesh Riemann–Stieltjes sense), where the limit is taken along any sequence of partitions $(\mathcal{P}^{n})_{n\in\mathbb{N}}$ of the interval $[s,t]$ such that $|\mathcal{P}^{n}|\to 0$ as $n\to\infty$, and comes with the estimate
@@ 2.2. Existence and uniqueness
Let $X,\tilde{X}\in D^{p}([0,T];\mathbb{R}^{d})$ for $p\in(2,3)$. For every $K>0$ there exist constants $C_{F}$ and $C_{F,K,X,\tilde{X}}>0$ such that the maps
@@
satisfy, for all $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}$, and every $0\leq s<r\leq t\leq T$,
@@
$\|F(Y)\|_{p,[s,t]}\leq C_{F}(1+\|Y\|_{p,[s,t]}),\quad\text{and}$
@@
$\|F(Y),F^{\prime}(Y,Y^{\prime});F(\tilde{Y}),F^{\prime}(\tilde{Y},\tilde{Y}^{\prime})\|_{X,p,[s,t]}$
@@
Moreover, there exists a componentwise non-decreasing function $K_{p}\colon[0,\infty)^{4}\to[0,\infty)$ such that
$\|Y,Y^{\prime}\|_{X,p}\leq K_{p}(|F_{0}(y)|,\|y,y^{\prime}\|_{X,p},C_{F},\|\mathbf{X}\|_{p}).$
@@
and introduce the subset of controlled paths
@@
$\quad\leq C_{F}(1+\|Y\|_{p,[0,t]})$
$\quad\lesssim_{p}C_{F}(1+(1+|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+\|R^{Y-y}\|_{\frac{p}{2},[0,t]}+\|R^{y}\|_{\frac{p}{2},[0,t]})$
$\quad\lesssim_{p}C_{F}(1+|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+3C_{F},$
@@
$\quad\lesssim_{p}C_{F}(1+\|Y,Y^{\prime}\|_{X,p,[0,t]})^{2}(1+\|X\|_{p,[0,t]})^{2}\|\mathbf{X}\|_{p,[0,t]}.$
Hence, for $t=t_{1}$ sufficiently small we obtain that $\mathcal{B}_{t_{1}}$ is invariant under $\mathcal{M}_{t_{1}}$. Note that $t_{1}$ depends only on $p$, $|F_{0}(y)|$, $C_{F}$, $\|\mathbf{X}\|_{p}$ as well as $\|y,y^{\prime}\|_{X,p}$.
Contraction. Let $(Y,Y^{\prime}),(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{B}_{t}$ for some $t\in(0,t_{1}]$. We have that
@@
where the implicit multiplicative constant depends only on $p$ and $\|\mathbf{X}\|_{p}$. Defining an equivalent norm on $\mathcal{V}^{p}_{X}$ by
$\|Y,Y^{\prime}\|^{(\delta)}_{X,p,[0,t]}:=\|Y^{\prime}\|_{p,[0,t]}+\delta\|R^{Y}\|_{\frac{p}{2},[0,t]},\quad\text{for }\delta\geq 1,$
@@
$\quad\lesssim C_{F,K,X,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t]}}{\delta}\bigg{)}\|Y-\tilde{Y},Y^{\prime}-\tilde{Y}^{\prime}\|^{(\delta)}_{X,p,[0,t]}.$
Hence, we can choose $\delta$ sufficiently large and $t=t_{2}\leq t_{1}$ sufficiently small such that
$C_{F,K,X,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t_{2}]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t_{2}]}}{\delta}\bigg{)}<1,$
where the left-hand side is up to a multiplicative constant which depends only on $p$ and $\|\mathbf{X}\|_{p}$. It follows that $\mathcal{M}_{t_{2}}$ is a contraction on the Banach space $(\mathcal{B}_{t_{2}},\|\cdot\|^{(\delta)}_{X,p,[0,t_{2}]})$. Hence, by Banach’s fixed point theorem, there exists a unique fixed point of the map  $\mathcal{M}_{t_{2}}$ , which is the unique solution of the RFDE (2.1) over the time interval $[0,t_{2}]$.
@@
We infer from Step 1 that there exists a constant $\gamma>0$, which depends only on $p$, $C_{F}$ and $C_{F,K,\mathbf{X},\mathbf{X}}$ such that, given any initial condition $\xi\in\mathcal{V}^{p}_{X}$, the local solution $(Y,Y^{\prime})$ established above exists on any interval $[s,t]$ such that $w(s,t)<\gamma$.
By [FZ18, Lemma 1.5], there exists a partition $\mathcal{P}=\{0=t_{0}<t_{1}<\dots<t_{N}=T\}$ of $[0,T]$, such that $w(t_{i},t_{i+1}-)<\gamma$ for every $i=0,1,\ldots,N-1$. We can then define the solution $(Y,Y^{\prime})$ on each of the half-intervals $[t_{i},t_{i+1})$. Given the solutions on $[t_{i},t_{i+1})$, the values $Y_{t_{i+1}}$ at the right end-point of the interval are uniquely determined by the jumps of $\mathbf{X}$ at time $t_{i+1}$. More precisely, let $y_{0;\cdot}=y_{\cdot}$, and define $y_{i}$, $i=1,\ldots,N-1$, by
$y_{i;t}=y_{t}+Y_{t_{i}-}-y_{t_{i}-}+F_{t_{i}-}(Y)X_{t_{i}-,t_{i}}+F^{\prime}_{t_{i}-}(Y,Y^{\prime})\mathbb{X}_{t_{i}-,t_{i}},\quad t\in[t_{i},t_{i+1}).$
Given the initial condition $(y_{i},y^{\prime}_{i})\in\mathcal{V}^{p}_{X}$, we obtain the solution $(Y,Y^{\prime})$ on $[t_{i},t_{i+1})$, $i=0,1,\ldots,N-1$. By pasting the solutions on each of these subintervals together, with $Y_{T}=y_{N;T}$, we obtain a unique global solution $(Y,Y^{\prime})$ of the RFDE (2.1) on the interval $[0,T]$.
Step 3: A priori estimate. It remains to show the existence of a componentwise non-decreasing function $K_{p}\colon[0,\infty)^{4}\to[0,\infty)$ such that
$\|Y,Y^{\prime}\|_{X,p}\leq K_{p}(|F_{0}(y)|,\|y,y^{\prime}\|_{X,p},C_{F},\|\mathbf{X}\|_{p}).$
From Step 2 and [FZ18, Lemma 1.5], we obtain the existence of a partition $\mathcal{P}=\{0=\tau_{0}<\tau_{1}<\ldots<\tau_{M}=T\}$ such that
(2.3) $\|Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})}\leq 4C_{F}+\|y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})}$
@@
for all $i=0,\ldots,M-1$. Here, $M$ depends only on $p$, $C_{F}$, $|F_{0}(y)|$, $\|y,y^{\prime}\|_{X,p}$, $\|\mathbf{X}\|_{p}$ and is, for $p$ fixed, non-decreasing in the other variables. Observe that
@@
it follows from (2.3) and (2.4) that
$\|Y,Y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1})}\leq|y^{\prime}_{\tau_{i}}|+|F_{\tau_{i}}(Y)|+4(1+C_{F})+\|y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})}+\|R^{y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1})}.$
Consequently, there exists a componentwise non-decreasing polynomial $Q_{p}^{(R)}\colon[0,\infty)^{4}\to[0,\infty)$ such that
$|R^{Y}_{\tau_{i+1}-,\tau_{i+1}}|\leq Q_{p}^{(R)}(|F_{\tau_{i}}(Y)|,\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$
@@
(2.5) $\|R^{Y}\|_{p,[\tau_{i},\tau_{i+1}]}\leq 1+\|R^{y}\|_{p,[\tau_{i},\tau_{i+1}]}+Q_{p}^{(R)}(|F_{\tau_{i}}(Y)|,\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$
for all $i=0,\ldots,M-1$. Moreover, since
$Y^{\prime}_{t-,t}=y^{\prime}_{t-,t}+F_{t-,t}(Y)X_{t-,t},$
@@
$\quad\leq\|y,y^{\prime}\|_{p,[\tau_{i},\tau_{i+1}]}+(|F_{\tau_{i}}(Y)|+C_{F}(1+\|Y\|_{p,[\tau_{i},\tau_{i+1})}))\|X\|_{p,[\tau_{i},\tau_{i+1}]}$
$\qquad+C_{F}(1+\|Y,Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})})^{2}(1+\|X\|_{p,[\tau_{i},\tau_{i+1}]})^{2}\|\mathbb{X}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}$

### v2

@@ 1. Introduction
The organization of the paper is as follows. In Section 2 we provide existence, uniqueness and continuity results for rough functional differential equations. In Section 3 we prove that various classes of rough differential equations are covered by the presented results on rough functional differential equations. In Section 4 we establish the existence of the Itô rough path lift of delayed martingales and discuss applications to stochastic differential equations with delay. Appendix A contains some auxiliary estimates for rough integrals.
@@ 2. Existence, uniqueness and continuity
Before treating rough functional differential equations (RFDEs), we recall the necessary definitions and some essentials from the theory of càdlàg rough paths, as introduced by Friz and Shekhar [FS17] and Friz and Zhang [FZ18]. The theory of càdlàg rough paths extends the classical rough path theory, allowing to deal with many stochastic processes with jumps [Che18], and has numerous applications, e.g., in probability theory [FZK23], numerical analysis [FZ18] and mathematical finance [ALP24].
@@ 2.1. Essentials on rough path theory
Throughout, let $T>0$ be a fixed finite time horizon. Let $\Delta_{T}:=\{(s,t)\in[0,T]^{2}\,:\,s\leq t\}$ be the standard $2$-simplex. A function $w\colon\Delta_{T}\to[0,\infty)$ is called a control function if it is superadditive, in the sense that $w(s,u)+w(u,t)\leq w(s,t)$ for all $0\leq s\leq u\leq t\leq T$. We write $w(s,t-):=\lim_{u\uparrow t}w(s,u)$ if $s<t$, and $w(s,t-):=0$ if $s=t$.
Whenever $(B,\|\cdot\|)$ is a normed space and $f,g\colon B\to\mathbb{R}$ are two functions on $B$, we shall write $f\lesssim g$ or $f\leq Cg$ to mean that there exists a constant $C>0$ such that $f(x)\leq Cg(x)$ for all $x\in B$. The constant $C$ may depend on the normed space, e.g. through its dimension or regularity parameters, and, if we want to emphasize the dependence of the constant $C$ on some particular variables, say $\alpha_{1},\ldots,\alpha_{n}$, then we will write $\lesssim_{\alpha_{1},\ldots,\alpha_{n}}$ or $C=C_{\alpha_{1},\ldots,\alpha_{n}}$. Unless otherwise stated, the dependence of the implicit constants on the variables is locally bounded; that is, if $\alpha_{1}\in A_{1},\ldots,\alpha_{n}\in A_{n}$, where $A_{1},\ldots,A_{n}$ are compact subsets of the range of $\alpha_{1},\ldots,\alpha_{n}$ respectively, then we have that $\sup_{\alpha_{1}\in A_{1},\ldots,\alpha_{n}\in A_{n}}C_{\alpha_{1},\ldots,\alpha_{n}}<\infty$.
@@
for $(s,t)\in\Delta_{T}$. We point out that, by definition, for $(s,t)\in\Delta_{T}$,
$|Y_{s,t}|\leq|Y^{\prime}_{s}||X_{s,t}|+|R^{Y}_{s,t}|\qquad\text{and}\qquad|Y^{\prime}_{t}|\leq|Y^{\prime}_{0}|+|Y^{\prime}_{0,t}|,$
which implies that
@@
exists (in the classical mesh Riemann–Stieltjes sense), where the limit is taken along any sequence of partitions $(\mathcal{P}^{n})_{n\in\mathbb{N}}$ of the interval $[s,t]$ such that $|\mathcal{P}^{n}|\to 0$ as $n\to\infty$. More precisely, in writing the product $Y_{u}X_{u,v}$, we apply the operator $Y_{u}\in\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k})$ onto $X_{u,v}\in\mathbb{R}^{d}$; and in writing the product $Y^{\prime}_{u}\mathbb{X}_{u,v}$, we use the natural identification of $\mathcal{L}(\mathbb{R}^{d};\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ with $\mathcal{L}(\mathbb{R}^{d}\otimes\mathbb{R}^{d};\mathbb{R}^{k})$. The rough integral comes with the estimate
@@
The estimate implies that $(\int_{0}^{\cdot}Y_{r}\,\mathrm{d}\mathbf{X}_{r},Y)\in\mathcal{V}_{X}^{p}([0,T];\mathbb{R}^{k})$ is a controlled path with respect to $X$, see also [FZ18, Remark 2.8].
@@
Let us now consider the rough functional differential equation (RFDE)
where $\mathbf{X}\in\mathcal{D}^{p}([0,T];\mathbb{R}^{d})$ is a càdlàg $p$-rough path for $p\in(2,3)$, $(y,y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$ is a given controlled path with respect to $X$ and further, where $(F,F^{\prime})\colon\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})\to\mathcal{V}^{p}_{X}([0,T];\mathcal{L}(\mathbb{R}^{d};\mathbb{R}^{k}))$ is a non-anticipative functional, i.e.
  • (i)
  • (ii)
for every $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}([0,T];\mathbb{R}^{k})$. The integral in (2.2) is defined as a (forward) rough integral, see (2.1) for its definition. Note that the RFDE (1.1) can be re-written in the form of (2.2), using a standard time-extension of the driving rough path.
@@ 2.2. Existence and uniqueness
For any $X,\tilde{X}\in D^{p}([0,T];\mathbb{R}^{d})$ and for every $K>0$, there exist constants $C_{F}>0$, which depends on $p$ and the functional $F$, and $C_{F,K,X,\tilde{X}}>0$, which additionally depends on $K$, $X$, $\tilde{X}$ such that the maps
@@
satisfy, for all $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, $(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{V}^{p}_{\tilde{X}}$, and every $0\leq s<t\leq T$,
@@
$|F_{t}(Y)|\leq C_{F},$
$|F_{t-,t}(Y)|\leq C_{F}(1+\|Y\|_{p,[s,t)}+|Y_{t-,t}|),$
$\|F(Y)\|_{p,[s,t]}\leq C_{F}(1+(|Y^{\prime}_{s}|+\|Y^{\prime}\|_{p,[s,t]})\|X\|_{p,[s,t]}+\|R^{Y}\|_{\frac{p}{2},[s,t]}),\quad\text{and}$
@@
$\|F(Y),F^{\prime}(Y,Y^{\prime});F(\tilde{Y}),F^{\prime}(\tilde{Y},\tilde{Y}^{\prime})\|_{X,\tilde{X},p,[s,t]}$
@@
Remark 2.2.
When refering to a functional $(F,F^{\prime})$ which satisfies Assumption 2.1, we consider a functional which maps a controlled path with respect to some path $X\in D^{p}$ to a controlled path with respect to $X$, where $X$ is not being fixed a priori. That is, $(F(Y),F^{\prime}(Y,Y^{\prime}))\in\mathcal{V}^{p}_{X}$ for $(Y,Y^{\prime})\in\mathcal{V}^{p}_{X}$, for any $X\in D^{p}$. Particularly, Assumption 2.1 is to hold for any $X,\tilde{X}\in D^{p}$, which are not assumed fixed.
@@
Moreover, there exists a componentwise non-decreasing function $K_{p}\colon[0,\infty)^{3}\to[0,\infty)$ such that
$\|Y,Y^{\prime}\|_{X,p}\leq K_{p}(\|y,y^{\prime}\|_{X,p},C_{F},\|\mathbf{X}\|_{p}).$
@@
noting that $\mathcal{M}_{t}(Y,Y^{\prime})$ is a controlled path with respect to $X$ as $\mathcal{V}^{p}_{X}$ is a Banach space, and introduce the subset of controlled paths
@@
$\quad\leq C_{F}(1+(|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+\|R^{Y}\|_{\frac{p}{2},[0,t]})$
$\quad\leq C_{F}+C_{F}(|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+C_{F}\|R^{Y-y}\|_{\frac{p}{2},[0,t]}+C_{F}\|R^{y}\|_{\frac{p}{2},[0,t]}$
$\quad\leq C_{F}(1+|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p,[0,t]})\|X\|_{p,[0,t]}+3C_{F},$
@@
$\quad\lesssim C_{F}(1+\|Y,Y^{\prime}\|_{X,p,[0,t]})^{2}(1+\|X\|_{p,[0,t]})^{2}\|\mathbf{X}\|_{p,[0,t]},$
where the implicit multiplicative constant depends only on $p$. Hence, for $t=t_{1}$ sufficiently small we obtain that $\mathcal{B}_{t_{1}}$ is invariant under $\mathcal{M}_{t_{1}}$. Note that $t_{1}$ depends on $p$, $|y^{\prime}_{0}|$, $C_{F}$ and $\|\mathbf{X}\|_{p}$.
Contraction. Let $(Y,Y^{\prime}),(\tilde{Y},\tilde{Y}^{\prime})\in\mathcal{B}_{t}$ for some $t\in(0,t_{1}]$, that is, setting $K:=5(1+\|y,y^{\prime}\|_{X,p}+C_{F})$, it holds that $\|Y,Y^{\prime}\|_{X,p,[0,t]},\|\tilde{Y},\tilde{Y}^{\prime}\|_{p,[0,t]}\leq K$. We have that
@@
where the implicit multiplicative constant depends on $p$ and $\|\mathbf{X}\|_{p}$. Defining an equivalent norm on $\mathcal{V}^{p}_{X}$ by
$\|Y,Y^{\prime}\|^{(\delta)}_{X,p,[0,t]}:=|Y^{\prime}_{0}|+\|Y^{\prime}\|_{p,[0,t]}+\delta\|R^{Y}\|_{\frac{p}{2},[0,t]},\quad\text{for }\delta\geq 1,$
@@
$\quad\lesssim C_{F,K,X,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t]}}{\delta}\bigg{)}\|Y-\tilde{Y},Y^{\prime}-\tilde{Y}^{\prime}\|^{(\delta)}_{X,p,[0,t]},$
where the implicit multiplicative constant depends on $p$ and $\|\mathbf{X}\|_{p}$. Hence, we can choose $\delta$ sufficiently large and $t=t_{2}\leq t_{1}$ sufficiently small such that
$C_{F,K,X,X}\bigg{(}(1+\delta)\|\mathbf{X}\|_{p,[0,t_{2}]}\vee\frac{1+\delta\|\mathbf{X}\|_{p,[0,t_{2}]}}{\delta}\bigg{)}\leq 1,$
where the left-hand side is up to a multiplicative constant which depends on $p$ and $\|\mathbf{X}\|_{p}$. It follows that $\mathcal{M}_{t_{2}}$ is a contraction on the subset of controlled paths $(\mathcal{B}_{t_{2}},\|\cdot\|^{(\delta)}_{X,p,[0,t_{2}]})$. Hence, by Banach’s fixed point theorem, there exists a unique fixed point of the map $\mathcal{M}_{t_{2}}$, which is the unique solution of the RFDE (2.2) over the time interval $[0,t_{2}]$.
@@
We infer from Step 1 that there exists a constant $\gamma>0$, which depends on $p$, $\|y,y^{\prime}\|_{X,p}$, $C_{F}$, $C_{F,K,X,X}$ and $\|\mathbf{X}\|_{p}$, such that the local solution $(Y,Y^{\prime})$ established above exists on any interval $[s,t]$ such that $w(s,t)\leq\gamma$, given any initial condition $\xi\in\mathcal{V}^{p}_{X}$ with $|\xi^{\prime}_{s}|\leq\|y,y^{\prime}\|_{X,p}$.
By [FZ18, Lemma 1.5], there exists a partition $\mathcal{P}=\{0=\tau_{0}<\tau_{1}<\dots<\tau_{N}=T\}$ of $[0,T]$, such that $w(\tau_{i},\tau_{i+1}-)<\gamma$ for every $i=0,1,\ldots,N-1$. We can then define the solution $(Y,Y^{\prime})$ on each of the half-intervals $[\tau_{i},\tau_{i+1})$. Given the solutions on $[\tau_{i},\tau_{i+1})$, the values $Y_{\tau_{i+1}}$ at the right end-point of the interval are uniquely determined by the jumps of $\mathbf{X}$ at time $\tau_{i+1}$. More precisely, let $y_{0;\cdot}=y_{\cdot}$, and define $y_{i}$, $i=1,\ldots,N-1$, by
$y_{i;t}=y_{t}+Y_{\tau_{i}-}-y_{\tau_{i}-}+F_{\tau_{i}-}(Y)X_{\tau_{i}-,\tau_{i}}+F^{\prime}_{\tau_{i}-}(Y,Y^{\prime})\mathbb{X}_{\tau_{i}-,\tau_{i}},\quad t\in[\tau_{i},\tau_{i+1}).$
We note that $|y^{\prime}_{i,\tau_{i}}|=|y^{\prime}_{i}|\leq\|y,y^{\prime}\|_{X,p}$. Given the initial condition $(y_{i},y^{\prime}_{i})\in\mathcal{V}^{p}_{X}$, we obtain the solution $(Y,Y^{\prime})$ on $[\tau_{i},\tau_{i+1})$, $i=0,1,\ldots,N-1$. By pasting the solutions on each of these subintervals together, with $Y_{T}=y_{N;T}$, we obtain a unique global solution $(Y,Y^{\prime})$ of the RFDE (2.2) on the interval $[0,T]$.
Step 3: A priori estimate. It remains to show the existence of a componentwise non-decreasing function $K_{p}\colon[0,\infty)^{3}\to[0,\infty)$ such that
$\|Y,Y^{\prime}\|_{X,p}\leq K_{p}(\|y,y^{\prime}\|_{X,p},C_{F},\|\mathbf{X}\|_{p}).$
As stated in Step 2, we obtain the existence of a partition $\mathcal{P}=\{0=\tau_{0}<\tau_{1}<\dots<\tau_{N}=T\}$ and define the solution on $(Y,Y^{\prime})$ on each of the half-intervals $[t_{i},t_{i+1})$, which then satisfies
(2.3) $\|Y^{\prime}\|_{p,[t_{i},t_{i+1})}\leq 4C_{F}+\|y^{\prime}\|_{p,[t_{i},t_{i+1})}$
@@
for all $i=0,\ldots,N-1$. Here, $N$ depends on $p$, $\|y,y^{\prime}\|_{X,p}$, $C_{F}$, $C_{F,K,X,X}$, $\|\mathbf{X}\|_{p}$ and is, for $p$ fixed, non-decreasing in the other variables. Observe that
@@
It follows from Assumption 2.1 (i), (2.3) and (2.4) that
$\|Y,Y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1})}\leq|y^{\prime}_{\tau_{i}}|+5(1+C_{F})+\|y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})}+\|R^{y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1})}.$
Consequently, there exists a componentwise non-decreasing polynomial $Q_{p}^{(R)}\colon[0,\infty)^{3}\to[0,\infty)$ such that
$|R^{Y}_{\tau_{i+1}-,\tau_{i+1}}|\leq Q_{p}^{(R)}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$
@@
(2.5) $\|R^{Y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}\leq 1+\|R^{y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}+Q_{p}^{(R)}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$
for all $i=0,\ldots,N-1$. Moreover, since
$Y^{\prime}_{t-,t}=y^{\prime}_{t-,t}+F_{t-,t}(Y),$
@@
$\quad\leq\|y,y^{\prime}\|_{p,[\tau_{i},\tau_{i+1}]}+(C_{F}+C_{F}(1+(|Y^{\prime}_{\tau_{i}}|+\|Y^{\prime}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1})})\|X\|_{p,[\tau_{i},\tau_{i+1}]}$
$\qquad+\|R^{Y}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1})})\|X\|_{p,[\tau_{i},\tau_{i+1}]}$
$\qquad+C_{F}(1+\|Y,Y^{\prime}\|_{p,[\tau_{i},\tau_{i+1})})^{2}(1+\|X\|_{p,[\tau_{i},\tau_{i+1}]})^{2}\|\mathbb{X}\|_{\frac{p}{2},[\tau_{i},\tau_{i+1}]}.$
Using (2.3) and (2.4), we can now conclude that there exist componentwise non-decreasing polynomials $Q_{p}^{(Y,J)},Q_{p}^{(Y^{\prime})}\colon[0,\infty)^{3}\to[0,\infty)$ such that
$|F_{\tau_{i+1}-,\tau_{i+1}}(Y)|\leq Q_{p}^{(Y,J)}(\|y,y^{\prime}\|_{X,p,[\tau_{i},\tau_{i+1}]},C_{F},\|\mathbf{X}\|_{p,[\tau_{i},\tau_{i+1}]})$